# Allowable Complex Black Holes in the Euclidean Gravitational Path Integral

### v1

@@ 3.4 KSW Conditions in the Bulk of Spacetime
In the plots collected in Figure 2, the regions where all the KSW conditions are satisfied are colored green. In our computations we found that the $p=0,1$ conditions are exhaustive, the $p\geq 2$ conditions do not anything new. This generalizes the analytical results near the asymptotic boundary presented in the previous section to the bulk or the geometry. In Figure 2, we observe the following features:
  • The KSW conditions become ever more stringent as one reads the Figure from top to bottom. This corresponds to motion from the horizon and the asymptotic boundary.
  • The strongest conditions in the parameter space are obtained at the asymptotic boundary (last row of plots).
@@
The spatial dependence of the instability is significant. That the breakdown is at large $r$ suggests that the black hole in Lorentzian signature becomes unstable in the region near the asymptotic boundary. That it is in a $2$-planes at $\vartheta=0,\pi/2$ indicates that the instability is driven by one component of the angular momentum. These features are consistent with a phase transition between the single center black hole and a two component configuration comprising both a black hole core and a BPS gas near the asymptotic boundary that carries large angular momentum efficiently. This two component configuration in Lorentzian spacetime is known as a Grey GalaxyKim et al. (2023); Bajaj et al. (2024).
@@ 4 Discussion
The divergence of the partition function (or index) in the grand canonical ensemble typically signals a breakdown of the chosen saddle point approximation and an instability towards a new phase. Our results suggest that the violation of the KSW conditions is the bulk geometric realization of this instability. Recently, the microcanonical version of the index has been explored in detail Choi et al. (2025a); Deddo et al. (2025); Jones et al. (2025), leading to the identification of novel two-component phases Kim et al. (2023); Choi et al. (2025b); Bajaj et al. (2024). These phases typically consist of a black hole at the center of the spacetime surrounded by a gas of gravitons (or a few dual giant gravitons). Notably, these multi-component phases dominate the ensemble in exactly the regions of parameter space where we found that the KSW criteria instruct us to exclude the standard complex Euclidean black hole saddles. This points to a compelling physical picture:
@@
  • Although these new two-component phases exist in a parameter space where the original saddle was unstable, the new configurations that one might construct to describe the endpoint of this instability should correspond to chemical potentials inside the convergent region of the index.
  • Crucially, these new phases were earlier Choi et al. (2025a) constructed in the Lorentzian signature, where the black hole component is extremal, supersymmetric, and real.
A fully Euclidean understanding of the novel two-component phases remains an open problem. The failure of the KSW criterion for the single-center black hole suggests a new, stable Euclidean saddle — perhaps a complex metric describing a “black hole plus gas” configuration — that effectively reproduces the features of microcanonical index.
Recently, Anand et al. (2025) (building on Bhattacharyya et al. (2008); Banerjee et al. (2012); Jensen et al. (2012); Shaghoulian (2016); Benjamin et al. (2024)) studied thermal partition functions near the boundaries of the convergence region. For 4D CFTs, they observed a semi-universal behavior as the chemical potentials $\Omega_{a}^{\prime}\to 0$ that is consistent with phase transitions between Kerr-AdS black holes, Grey Galaxies and pure Thermal AdS gas. It would be highly desirable to see an explicit phase transition in the superconformal index in the canonical ensemble that mirrors the Grey Galaxy phase transitions seen in the thermal partition function. The numerical evidence from the microcanonical ensemble Choi et al. (2025a); Deddo et al. (2025) strongly supports the existence of supersymmetric versions of the Grey Galaxies, but an analytic Euclidean solution describing them is still lacking.
Finally, while this work focused on the KSW criterion, other proposals for determining the admissibility of complex metrics in the gravitational path integral have been explored recently Di Tucci et al. (2020); Jonas et al. (2022); Mahajan and Singhi (2025); Singhi (2025). It would be instructive to apply these alternative diagnostics to the unequal angular momenta saddles studied here. Specifically, if those results can reproduce the same forbidden regions as the KSW criterion, it would provide progress towards a universal rule governing admissibility of complex metrics in Euclidean quantum gravity.

### v3

@@ 3.4 KSW Conditions in the Bulk of Spacetime
In the plots collected in Figure 2, the regions where all the KSW conditions are satisfied are colored green. In our computations we found that the $p=0,1$ conditions are exhaustive, the $p\geq 2$ conditions do not add anything new. This generalizes the analytical results near the asymptotic boundary presented in the previous section to the bulk of the geometry. In Figure 2, we observe the following features:
  • The KSW conditions become ever more stringent as one reads the Figure from top to bottom. This corresponds to motion from the horizon and the asymptotic boundary. The strongest conditions are obtained at the asymptotic boundary, shown in the last row of plots.
@@
The spatial dependence of the instability is significant. That the breakdown is at large $r$ suggests that the black hole in Lorentzian signature becomes unstable in the region near the asymptotic boundary. That it is in a $2$-planes at $\vartheta=0,\pi/2$ indicates that the instability is driven by one component of the angular momentum. These features are consistent with a phase transition between the single center black hole and a two component configuration comprising both a black hole core and a BPS gas near the asymptotic boundary that carries large angular momentum efficiently. Such a two component configuration in Lorentzian spacetime is known as a Grey GalaxyKim et al. (2023); Bajaj et al. (2024).
@@ 4 Discussion
The divergence of the partition function (or index) in the grand canonical ensemble typically signals a breakdown of the saddle point approximation and an instability towards a new phase. Our results suggest that the violation of the KSW conditions is the bulk geometric realization of this instability. Recently, the microcanonical version of the index has been explored in detail Choi et al. (2025a); Deddo et al. (2025); Jones et al. (2025), leading to the identification of novel two-component phases Kim et al. (2023); Choi et al. (2025b); Bajaj et al. (2024). These phases typically consist of a black hole at the center of the spacetime surrounded by a gas of gravitons (or a few dual giant gravitons). Notably, these multi-component phases dominate the ensemble in exactly the regions of parameter space where we found that the KSW criteria instruct us to exclude the standard complex Euclidean black hole saddles. This points to a compelling physical picture:
@@
  • Although these new two-component phases exist in a parameter space where the original saddle was unstable, the new two component configurations that describe the endpoint of this instability must have chemical potentials inside the convergent region of the index.
  • These new phases were earlier constructed in the Lorentzian signature, where the black hole component is extremal, supersymmetric, and real Choi et al. (2025a) .
A fully Euclidean understanding of the novel two-component phases remains an open problem. The failure of the KSW criterion for the single-center black hole suggests a new, stable Euclidean saddle — perhaps a complex metric describing a “black hole plus gas” configuration — that reproduces the features of the microcanonical index.
Recently, Anand et al. (2025) (building on Bhattacharyya et al. (2008); Banerjee et al. (2012); Jensen et al. (2012); Shaghoulian (2016); Benjamin et al. (2024)) studied thermal partition functions near the boundaries of the convergence region. For 4D CFTs, they observed a semi-universal behavior as the chemical potentials $\Omega_{a}^{\prime}\to 0$ that is consistent with phase transitions between Kerr-AdS black holes, Grey Galaxies and the pure Thermal AdS gas. It would be highly desirable to see an explicit phase transition in the superconformal index in the canonical ensemble that mirrors the Grey Galaxy phase transitions seen in the thermal partition function. The numerical evidence from the microcanonical ensemble Choi et al. (2025a); Deddo et al. (2025) strongly supports the existence of supersymmetric versions of the Grey Galaxies, but Euclidean saddles describing them are still lacking.
Finally, while this work focused on the KSW criterion, other proposals for determining the admissibility of complex metrics in the gravitational path integral have been explored recently Di Tucci et al. (2020); Jonas et al. (2022); Mahajan and Singhi (2025); Singhi (2025); Ailiga et al. (2026). It would be instructive to apply these alternative diagnostics to the unequal angular momenta saddles studied here. Specifically, if those results can reproduce the same forbidden regions as the KSW criterion, it would provide progress towards a universal rule governing admissibility of complex metrics in Euclidean quantum gravity.